How many moles are there in:






(a) 6g of Magnesium





(b) 88g of CO2

1
Expert's answer
2022-02-04T09:42:03-0500

a)Moles= mass/ rfm

Moles= 6/24.3= 0.247moles

b) moles= 88/44= 2.0moles